OpenStudy (anonymous):

Solve the homogeneous differential equation 2x dx + (y-3x)dy = 0

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

Have you done these yet? Bạn đã làm những chưa?

OpenStudy (loser66):

roi

OpenStudy (loser66):

hahahaha.... interesting. !!!

OpenStudy (anonymous):

does that mean no? lol i don't speak vietnamese i just have a translator

OpenStudy (loser66):

yes! I did

OpenStudy (loser66):

how can you put it there?

OpenStudy (anonymous):

copy and paste from translate.google.com

OpenStudy (loser66):

you are...but interesting

OpenStudy (anonymous):

so how do you solve @Loser66 there's two approaches to homogenous de's and how do you know if it is homogeneous? vậy làm thế nào để bạn giải quyết @ Loser66 có hai phương pháp tiếp cận đồng nhất de và làm thế nào để bạn biết nếu nó là đồng nhất?

OpenStudy (loser66):

hahahaah... I know how to consider whether it's exact or not, friend. there are 2 ways to solve.separate or substitute y =Vx

OpenStudy (loser66):

am I right?

OpenStudy (anonymous):

yes you substitute y=vx... the way to tell is a bit hard to explain, but i'll try

OpenStudy (anonymous):

so say the equation is of the form M(x,y)dx + N(x,y)dy If you rewrite the equation as M(xt,yt)dx + N(xt,yt)dy and the same factor of t factors out of the equation, then its homogenous... does that make sense?

OpenStudy (anonymous):

So M(xt,yt) = 2xt = t(2x) and N(xt,yt) = yt -3xt = t(y-3t) both factor out a t so its homogenous...

OpenStudy (loser66):

I have a way to figure out the solution: let see.

OpenStudy (loser66):

2xy -3x^2 =C, it's mine

OpenStudy (anonymous):

I got something way different

OpenStudy (loser66):

ok, that's the trick. let ,me redo, heheheh

OpenStudy (anonymous):

i didn't check if that works, (i hate the calculus of checking it in this form heh)

OpenStudy (anonymous):

2x dx +(y - 3x) dy = 0 let y = ux . So dy = u dx + x du and use substitution. 2x dx + (ux - 3x) ( u dx + x du) = 0 2 dx + (u - 3)(u dx + x du) = 0 2 dx + u^2 dx + ux du - 3u dx - 3x du = 0 (u^2 -3u +2)dx = ((3x -ux) du (u-2)((u-1)dx = x (3 - u) du now separate the variables. dx/x = (3-u) / (u-2) (u -1) du the left hand side is ln x + C. The right hand side will require partial fractions. It will be of the form A/ (u-2) + B/ (u-1). Solving for A and B you will get A = 1 and B = -2 . So integrating the right hand side means integrating 1/(u-2) + -2/(u-1). We obtain ln(u-2) -2ln(u-1). We now have: lnx + C = ln(u-2) -2ln(u-1) Substitute back in that u = y/x (from the beginning substitution) We now have: lnx + C = ln((y-2x)/x) -2ln((y-x)/x) Apply log properties: lnx + C = ln(y-2x) - ln x -2ln(y-x) + 2ln x C = ln(y-2x) -2ln(y-x)

OpenStudy (anonymous):

\[\frac{ y^2-2xy+x^2 }{ y-2x }=C\] is the answer but i was looking to see if you did it differently @loser66

OpenStudy (anonymous):

@rrr same answer as mine, except you need to raise both sides to e to cancel the logs

OpenStudy (anonymous):

our answers are reciprocals of each other though, hrmm

OpenStudy (loser66):

to me, I check whether it is exact or not. yours is not. then find \[\mu\] to make it exact. by formula. base on that differential equation which has exact solution find int respect to y of M and int respect to x of N. pick like term add them together to get the solution.

OpenStudy (anonymous):

It is OK that they are reciprocals since they would equal an arbitrary constant. The constants that they equal are not the same constant. They are also reciprocals.

OpenStudy (anonymous):

Let me make sure I understand your question. You want to know what on your particular solution, you have a t in one solution and not in the other. when you are making it look like your RHS of the nonhomogeneos DE.

OpenStudy (loser66):

this is another one

OpenStudy (anonymous):

When you find your from of particular solution when you working with the method of undetermined coefficient, your form is determined not only by the RHS of your DE but also by the complementary solution of the DE.

OpenStudy (loser66):

oh yes, I got what you mean, because the solution from the LHS has the form of e^-t already, so, the nonhomogeneous part must be At^-t , right? On example 1, the LHS doesn't have the form of e^2t so no need to have t at the RHS, right?

OpenStudy (anonymous):

Yes, you have it.

OpenStudy (loser66):

the last one, please, please, please!!

OpenStudy (anonymous):

Gotta go put my kids to bed. Sorry.

OpenStudy (anonymous):

what's the question?

OpenStudy (loser66):

at my attachment above

OpenStudy (anonymous):

\[y"-3y'-4y=-8e^tcos(t)\] this?

OpenStudy (loser66):

yes.

OpenStudy (anonymous):

so we need the particular solution

OpenStudy (loser66):

cos (2t)

OpenStudy (loser66):

\[\huge y"-3y'-4y=-8e^tcos(2t)\]

OpenStudy (anonymous):

ok got it, give me a sec ok?

OpenStudy (loser66):

sure. I am waiting

OpenStudy (anonymous):

been just a bit since i've done nonhomogenous 2nd order equations, and there's a bunch of different ways to do this

OpenStudy (loser66):

I need just one. give me it, please,

OpenStudy (anonymous):

\[y_p=e^t(Acos(2t)+Bcos(2t)\] is what you used yes?

OpenStudy (anonymous):

lol except the second is sin(2t)

OpenStudy (loser66):

I copied from my prof

OpenStudy (loser66):

yes, you are wrong there. but no big deal

OpenStudy (anonymous):

ok so you know that the right hand side is going to be of that form

OpenStudy (loser66):

eliminate one constant in the front of e^t , right?

OpenStudy (loser66):

then?

OpenStudy (anonymous):

yes we need a constant infront to... ok now we need to find the first and second derivatives and plug this all in to the ODE

OpenStudy (anonymous):

which is not easy, can make mistakes

OpenStudy (loser66):

\[e^t(Acos2t+Bsin2t)=De^{(1+2i)t}\] right? from Euler's formula for complex number, right?

OpenStudy (anonymous):

yes you can do that too

OpenStudy (loser66):

Oooookk, I got it now, thanks a ton .

OpenStudy (anonymous):

ok :)

OpenStudy (loser66):

hey, you did something on justin.tv for fun?

OpenStudy (anonymous):

Did i see something for fun? there's all kinds of things on justin.tv

OpenStudy (anonymous):

oh no my channel on justin.tv isn't just for fun, its part of my business

OpenStudy (loser66):

what are you doing there?

OpenStudy (anonymous):

hang on a second, i'm helping someone else